# Shahriar Noroozizadeh

### v2

@@ 2.2 Success of implicit in-weights reasoning
Our first observation is that, despite the adversarially-constructed topology of our graph, and even on path lengths of $6$ to $10$ hops and graphs as large as $10^{4}$ nodes, our next-token-trained models succeed at the in-weights path-finding task. We report this in Figs.˜4 and 14 (left). This isolates an implicit in-weights reasoning sharper than what exists in the literature. Known instances are on small scales of 200 or fewer entities [85, 48] or on 2-hop tasks [174, 43, 189, 74]. We report success in some harder topologies in §C.2.
Next, we isolate an even sharper instance of success. To arrive at that, we ask: why does the next-token learner succeed on massive in-weights path-star tasks, when back in B&N’24 it failed on even tiny in-context path-star tasks? A possible factor could be that in the in-weights setting, the model does not succumb to trivial left-to-right, next-token shortcuts; thus gradients from farther-away tokens survive, allowing the model to discover the true right-to-left solution. However, we find this is likely not the case since the tokens are not learned in a right-to-left order—we report this in Figs.˜4 and 14 (middle) and elaborate on it in §A.1.1. This inspires us to test a harder task where any future-token supervision is manually eliminated: we train the model purely on the first-token loss, conditioned on only the leaf node that is many hops away. Learning the first token—the key decision-making token—in isolation, is (theoretically and empirically) impossible in the in-context path-star task, as B&N’24 show and as reproduced in Fig.˜5 (right). In striking contrast in our in-weights task, our model learns this token in isolation, demonstrating an even sharper instance of implicit reasoning:
(Hardest, first token is learned in isolation) On in-weights path-star graphs of as many as $5\times 10^{4}$ nodes, when the model is trained on only edge-memorization examples and first-token-training examples from $75\%$ of all the paths, on unseen paths, both the Transformer and Mamba are able to predict the first token when conditioned on held-out leaves, with as much as 100% accuracy (Figs.˜4 and 14 (right)).
That the hardest (first) token is learned, we argue, is difficult to square with the abstraction that parametric memory strictly stores local associations. We explain this in the next subsection.
Figure 4: Success of Transformer in in-weights path-star task. (§2) (left) A next-token-trained Transformer achieves perfect or highly non-trivial accuracy on large path-star graphs $\mathcal{G}_{d,\ell}$. (middle) Learning order of tokens. The tokens of a path are not learned in the reverse order i.e., the model does not learn the right-to-left solution. Thus, gradients from the future tokens are not critical for success. (right) Success of hardest-token-only task. In fact, the hardest token (the first) given the leaf is learned in isolation to non-trivial accuracy (Section˜2.2). Success of this $\ell$-fold composition task is hard to explain within the associative memory view (§2.3). Analogous plots for Mamba are in Fig.˜14. We invite the reader to contrast these in-weights task results with the in-context task ones in Fig.˜5.
Figure 5: Failure of Transformer in in-context path-star task. (B&N’24) We report the failure of next-token Transformers in the in-context version of the path-star task, reproducing results from B&N’24. (left) Full path accuracy remains at chance level across different small graph sizes. (middle) Learning order of tokens with teacherless (multi-token-trained) objective shows a clear right-to-left learning cascade. (right) The hardest (first) token given the leaf fails to be learned in isolation, contrasting sharply with in-weights success shown in Fig.˜4.
@@ 2.3 The contradiction behind learning the hardest token
Atomic facts are often imagined to be stored in the parameters as local associations, where a matrix acts as a lookup table. We make this concrete. For any tokens $u$ and $v$, let $f(u)[v]$ denote the logit of a sequence model $f$ on the target $v$, conditioned on the input $u$. Then:
@@
(Local associative parametric memory) We say that a deep sequence model $f$ has memorized a graph $\mathcal{G}=({V},E)$ associatively iff for any vertices $u$ and $v$, the logit $f(u)[v]$ is high only on adjacent vertices, $(u,v)\in E$. In its simplest form, this model is typically abstracted as $f(u)[v]={\bf{\Phi}}(u){\bf{W}}_{\tt assoc}{\bf{\Phi}}(v)$. Here, ${\bf{W}}_{\tt assoc}$ encodes the adjacencies as $\sum_{(u,v)\in E}{\bf{\Phi}}(u)\otimes{\bf{\Phi}}(v)$; the embeddings $\bf{\Phi}$} are arbitrary “keys” that by themselves encode no associations in graph $\mathcal{G}$, e.g., they are orthogonal embeddings [177, 51, 78, 202, 27] or random (and thus, near-orthogonal) embeddings [124, 26, 20, 176].
@@
Observe that, in this abstraction, the first node in our task requires composing a local recall function $\ell$-many times: recall the predecessor of ${v}_{\tt{leaf}}$ from the weights, namely $v_{\ell-1},$ then the predecessor of that, $v_{\ell-2}$, and so on. In short, one can write $v_{1}=\texttt{Predec}^{\ell}({v}_{\tt{leaf}})$.
@@
Hardness of learning compositions. With this structure however, the first-token $\ell$-fold composition task should intuitively require $\Omega(\exp(\ell))$ compute to learn—just like the first token was demonstrably hard-to-learn in the in-context path-star task [B&N’24]. One intuition is that the optimization task is a search for a needle in the haystack: there is an $\exp(\ell)$ space of possible discrete compositions, and all but the correct one have an equally miserable loss value; with the loss terrain rendered flat and the gradients uninformative, the learner is forced to sift through the vast hypothesis space to find a needle. As a preliminary corroboration of this intuition, in §E.1, we find that models fail at this task when the embeddings are frozen. We also refer the reader to existing results that empirically demonstrate that compositions are hard to learn [156, 2, 61, 54, 1, 2] and others that theoretically demonstrate hardness [108, 32, 132, 156, 180, 159] (although only in a certain sense; see Remark˜2).
@@
The power of step-wise aid. This barrier could be surmounted if, rather than providing supervision from only the end output of the composition, there was supervision from each hop. This could mean providing a data curriculum [149, 180, 188, 5, 36] or providing chain-of-thought supervision [181, 125] or reducing compositionality through various means. Indeed, prior positive results on implicit in-weights reasoning involve one such aid or the other. First, Khona et al. [85] have paths of varying lengths (see their Fig 11), which provides an implicit curriculum; they also provide full path supervision, not just the first-token loss. Furthermore, their test and train paths overlap, reducing compositionality, potentially allowing the model to stitch substrings it has seen (an ability demonstrated as possible [177]). Other results [174, 43, 189, 74] are on 2-hop tasks, implying a friendly exponent.
@@
Our setting, by design, offers none of these aids. Our paths have fixed lengths (hence, no implicit curriculum), are disjoint (so, no test-train overlaps), and require as much as a $10$-fold composition (so, a daunting exponent). What we isolate, therefore, is a stronger and more sterile instance of implicit in-weights reasoning, one that is cleanly inconsistent within the associative view of parametric memory.
@@ 2.4 Two competing data structures for parametric memory
The paradox begins to resolve with fragmented yet salient observations in literature—[78, 85, 189] and [74]—that we will flesh out more generally. First, [78, 85, 189] find that the embeddings of their tokens reflect some useful notion of distance, aiding their respective implicit reasoning tasks. Independently, [74, 152] find that memorizing associations through an explicitly factorized architecture allows the model to draw unseen (two-hop) inferences. We will generalize these observations to our much larger graphs and many-hop tasks; but first, we point out that these observations are related and lead to profound implications. First, these observations paint an altogether different picture of parametric memory itself: a memory of atomic facts that does not take shape as a lookup over arbitrary embeddings but as a geometry of highly-organized embeddings. While the former directly stores the adjacency matrix $\bf{A}$} in the matrix ${\bf{W}}_{\tt assoc}$ (up to some rotations defined by $\bf{\Phi}$}), the latter “factorizes” the associations into a low-rank form, ${\bf{\Phi}}_{\tt geom}(u)^{T}{\bf{\Phi}}_{\tt geom}(v)$. Importantly, whereas an associative memory only makes local information accessible, a geometric memory readily exposes global multi-hop relationships (§4 discusses what these global relationships are). We define the notion of geometric memory below, and compare the two forms of memory in Table˜1.
(a) Edge-Memorized and Full-Path-Trained
(b) Model Embeddings
(c) Only Edge-Memorized
Figure 6: Evidence of global geometry of Transformer in path-star task. (a) In the heatmap, entry $(i,j)$ is the cosine distance between the leaf embedding of path $i$ (row) and the first-hop embedding of path $j$ (col). The clear diagonal line implies that embeddings within each path are more aligned, reflecting global structure. (b) UMAP projection of token embeddings where each point is a node embedding; color indicates path identity. Different paths form separated clusters (see Fig.˜27 for clearer image). (c) For a later experiment in §3.1: heatmap for a model trained only on edge-memorization still reveals a level of geometry, although weaker than in (a). Analogous plots for Mamba are in Fig.˜15
(Global geometric parametric memory) We say that a deep sequence model $f$ has memorized a graph $\mathcal{G}=({V},E)$ geometrically iff for any pairs of vertices $u$ and $v$, the logit $f(u)[v]$ reflects some notion of multi-hop closeness in $\mathcal{G}$. At its simplest, this can be abstracted as $f(u)[v]={\bf{\Phi}}_{\tt geom}(u)\cdot{\bf{\Phi}}_{\tt geom}(v)$, where ${\bf{\Phi}}_{\tt geom}$ reflects some graph embedding of the vertices. [^1] In the simple settings of this paper, the geometric embeddings ${\bf{\Phi}}_{\tt geom}$ in practice are found in the token embeddings themselves; but more generally, such an embedding may reside in any other layer. We consider even that to be geometric memorization. [^2] Note that this simplistic abstraction may assign the highest output logit to the input vertex itself (i.e., $\arg\max_{v}{\bf{\Phi}}_{\tt geom}(u)\cdot{\bf{\Phi}}_{\tt geom}(v)=u$) regardless of whether a self-loop $(u,u)$ exists. To account for this, one may incorporate an auxiliary self-loop-detecting circuit. We treat this as a pedantic detail best ignored for a tidy discussion.
@@
The resolution. One can view the two parametric memories of Sections˜2.3 and 2.4 as two competing data structures, both representable by a deep sequence model, each yielding its own learning complexity for the hardest token (much like how a heap or an array would yield different search complexities). The local associative memory is analogous to a linked list: for the first token, this would incur a (benign) $\ell$-hop lookahead step but learning this incurs an exponential cost. The geometric memory is powerful in that it reduces this learning complexity all the way to $\Theta(1)$. For instance, we may expect that the embeddings of all nodes in path $i$ are clustered tightly around a unique path vector $\mathbf{z}_{i}$. In this data structure, learning the first token given the leaf node is a one-hop task: simply find a one-hop neighbor of the central node that is best-aligned with ${\bf{\Phi}}_{\tt geom}({v}_{\tt{leaf}}^{(i)})$. As stated below, such a structure indeed materializes both in our massive graphs and in tiny graphs; this geometry turns out to be fundamental to any gradient-descent-trained deep sequence model and not specific to attention or residual connections:
@@
(Evidence of global geometry across various deep sequence models) For our large path-star graphs, the (token) embeddings of the leaf and first node of a path are clustered closer to each other than those of other paths in both the Transformer and Mamba (Fig.˜6 and Fig.˜15). Similar geometries arise for a variety of small graphs for the Transformer, Mamba and a simple 3-layer neural network model without attention and residual connections—see visualizations in Fig.˜1 and Section˜C.3; §C.3: Figs.˜18, 19, 20 and 21) and node-node cosine-similarity heatmaps (§C.3; Figs.˜24 and 23).
@@
While these geometries help make sense of why implicit reasoning succeeds, it also leaves us with a foundational question of why the model prefers one form of memory over the other. We elaborate on this in the next section.
@@ 3 The memorization puzzle: The emergence of geometric memory is not easy to explain
The reader may offer various plausible explanations for this behavior by borrowing intuition from other well-known phenomena. Some readers may find the rise of a geometry familiar in light of well-known geometries of features in language [40, 130, 112] and arithmetic tasks [100, 120, 117, 59]. Some others may find the phenomenon reminiscent of the competition between memorization and generalization studied in the generalization puzzle [195, 122, 17, 23] in regression and classification settings. These settings suggest that representations arise due to pressures from either the supervision or the architecture or the optimizer; such pressures forbid a lookup. Thus, it may be tempting to cite these well-known learning pressures as plausible explanations for what we observe in our sequence models. However, we design experiments refuting such explanations: we show that a geometric representation arises over the associative lookup even in the absence of such learning pressures. Through this analysis, we pose the memorization puzzle in sequence modeling.
@@ 3.1 Supervisory pressure does not explain geometric memory
Geometric representations arise in various tasks with reasoning objectives, from arithmetic [100, 120, 117, 59] to path-finding [85] to two-hop reasoning [174, 43, 189, 74]. Reasoning supervision has also been shown to act as a regularizer that elicits useful structures in the model’s storage [74]. This trail of findings overwhelmingly directs us to the following guess for why we see a geometry:
(Pressure from global supervision) Perhaps, the model memorizes via a global geometry rather than via local associations due to the supervisory pressure of a multi-hop reasoning objective.
This hypothesis however is weakened by the distinctive design of the path-star task. First, the test paths are untouched by any path-finding supervision, yet these paths exhibit a global geometry. Next, in the hardest-token-only task, path-finding supervision is applied only on the path end-points; yet a geometry materializes on intermediate nodes (see Fig.˜25(b)). As an even cleaner refutation of this hypothesis, we ablate the reasoning supervision, and train purely on local supervision i.e., on edge memorization:
(Geometry arises even from memorizing local, atomic co-occurrences) A global geometry emerges even in locally-supervised models for various architectures, as seen in the heatmaps for the large path-star graph (Figs.˜6(c) and 15(c)). A locally-supervised model can be subsequently finetuned purely on the hardest-token task and achieve high test accuracy on unseen paths (§D.3). We also note that the geometries seen on the small graphs (Fig.˜1; §C.3) are for various architectures trained only on edge-memorization.
From this, we isolate the following insight: geometries can arise instead of an associative lookup, even when memorizing atomic co-occurrences devoid of any pressure to reason or to find high-level, multi-hop structures in data.
3.2 Capacity or optimizer pressures do not explain geometric memory.
We may seek yet other explanations out of canonical intuitions in theories of generalization. In regression or classification, conventional wisdom was that models learn representations rather than a lookup when there is an explicit capacity pressure e.g., the architecture or the regularizers forbid a lookup. In modern theories since the popularization of the generalization puzzle [122, 195], renewed wisdom has been that this pressure is implicit: even if the model can express a lookup, gradient descent implicitly prefers solutions that are easier-to-find or are more succinct, thus leading to useful representations. It may seem that one of these theories, when imported to the world of sequence models, can tell us why a geometric representation arises over a naive lookup of atomic associations.

### v3

@@ 2.2 Success of implicit in-weights reasoning
The path-star graph is adversarially constructed and involves a particularly hard-to-learn token. Yet, in our in-weights version of this task, we find that next-token-trained models succeed at path-finding, even on path lengths of $6$ to $10$ hops and graphs as large as $10^{4}$ nodes. In fact, even when we manually eliminate all supervision from the later tokens in the path, and train only on the loss of the first token loss—the key decision-making token—the model predicts this token successfully on unseen paths (Fig.˜3). This is striking because learning the first token in isolation, one would intuitively expect, is (theoretically and empirically) hard—an intuition we will shortly elaborate on. This success also isolates an instance of implicit in-weights reasoning sharper than what exists in the literature. Known instances are on small scales of 200 or fewer entities [91; 53] or on 2-hop tasks [184; 47; 200; 80]. We report success on some harder topologies in §C.2.
(Hardest, first token is learned in isolation) On in-weights path-star graphs of as many as $5\times 10^{4}$ nodes, when the model is trained on only edge-memorization examples and first-token-training examples from $75\%$ of all the paths, on unseen paths, both the Transformer and Mamba are able to predict the first token when conditioned on held-out leaves, with as much as 100% accuracy (Figs.˜3 and 14 (right)).
@@ 2.3 The contradiction behind learning the hardest token
That the hardest (first) token is learned, we argue, is difficult to square with the default abstraction of memory in neural networks. Atomic facts are often imagined to be stored in the parameters as local associations, where a matrix acts as a lookup table. We make this concrete. For any tokens $u$ and $v$, let $f(u)[v]$ denote the logit of a sequence model $f$ on the target $v$, conditioned on the input $u$. Then:
@@
(Local associative parametric memory) We say that a deep sequence model $f$ has memorized a graph $\mathcal{G}=({V},E)$ associatively iff for any vertices $u$ and $v$, the logit $f(u)[v]$ is high only on adjacent vertices, $(u,v)\in E$. In its simplest form, this model is typically abstracted as $f(u)[v]={\bf{\Phi}}(u)^{T}{\bf{W}}_{\tt assoc}{\bf{\Phi}}(v)$. Here, ${\bf{W}}_{\tt assoc}$ encodes the adjacencies as $\sum_{(u,v)\in E}{\bf{\Phi}}(u)\otimes{\bf{\Phi}}(v)$; the embeddings $\bf{\Phi}$} are arbitrary “keys” that by themselves encode no associations in graph $\mathcal{G}$, e.g., they are orthogonal embeddings [187; 56; 84; 215; 29] or random (and thus, near-orthogonal) embeddings [133; 28; 22; 186].
@@
Observe that, in this abstraction, the first node in our task requires implicitly composing a local recall function $\ell$-many times: (implicitly) recall the predecessor of ${v}_{\tt{leaf}}$ from the weights, namely $v_{\ell-1},$ then the predecessor of that, $v_{\ell-2}$, and so on (all without an explicit chain of thought). In short, one can write $v_{1}=\texttt{Predec}^{\ell}({v}_{\tt{leaf}})$.
@@
Hardness of learning compositions. With this structure however, the first-token $\ell$-fold composition task should intuitively require $\Omega(\exp(\ell))$ compute to learn—just like the first token was demonstrably hard-to-learn in the in-context path-star task [B&N’24]. One intuition is that the optimization task is a search for a needle in the haystack: there is an $\exp(\ell)$ space of possible discrete compositions, and all but the correct one have an equally miserable loss value; with the loss terrain rendered flat and the gradients uninformative, the learner is forced to sift through the vast hypothesis space to find a needle. As a preliminary corroboration of this intuition, in §E.1, we find that models fail at this task when the embeddings are frozen. We also refer the reader to existing results that empirically demonstrate that compositions are hard to learn [166; 2; 67; 59; 1; 2] and others that theoretically demonstrate hardness [117; 35; 142; 166; 191; 169] (although only in a certain sense; see Remark˜2).
@@
The power of step-wise aid. This compositionality barrier could be surmounted if, rather than providing supervision from only the end output of the composition, there was supervision from each hop. This could mean providing a data curriculum [159; 191; 199; 5; 40] or providing chain-of-thought supervision [192; 135] or reducing compositionality through various means. Indeed, prior positive results on implicit in-weights reasoning involve one such aid or the other. First, Khona et al. [91] have paths of varying lengths (see their Fig. 11), which provides an implicit curriculum; they also provide full path supervision, not just the first-token loss. Furthermore, their test and train paths overlap, reducing compositionality, potentially allowing the model to stitch substrings it has seen (an ability demonstrated as possible [187]). Other results [184; 47; 200; 80] are on 2-hop tasks, implying a friendly exponent.
@@
Our setting, by design, offers none of these aids. Our paths have fixed lengths (hence, no implicit curriculum), all later token gradients eliminated (so no supervision from the intermediate hops), the paths disjoint (so, no test-train overlaps), and there is as much as a $10$-fold composition (so, a daunting exponent). What we isolate, therefore, is a stronger and more sterile instance of implicit in-weights reasoning, one that is cleanly inconsistent within the associative view of parametric memory.
@@ 2.4 Two competing data structures for parametric memory
The paradox begins to resolve with fragmented yet salient observations in literature—[84; 91; 200] and [80]—that we will flesh out more generally. First, Jiang et al. [84]; Khona et al. [91]; Ye et al. [200] find that the embeddings of their tokens reflect some useful notion of distance, aiding their respective implicit reasoning tasks. Independently, Huang et al. [80]; Saxe et al. [162] find that memorizing associations through an explicitly factorized architecture allows the model to draw unseen (two-hop) inferences. We will generalize these observations to our much larger graphs and many-hop tasks; but first, we point out that these observations are related and lead to profound implications. First, these observations paint an altogether different picture of parametric memory itself: a memory of atomic facts that does not take shape as a lookup over arbitrary embeddings but as a geometry of highly-organized embeddings. While the former directly stores the adjacency matrix $\bf{A}$} in the matrix ${\bf{W}}_{\tt assoc}$ (up to some rotations defined by $\bf{\Phi}$}), the latter factorizes the associations into ${\bf{\Phi}}_{\tt geom}(u)^{T}{\bf{\Phi}}_{\tt geom}(v)$. Importantly, whereas an associative memory only makes local information accessible, a geometric memory readily exposes global multi-hop relationships. We define the notion of geometric memory below, and compare the two forms of memory in Table˜1. (The definition below intentionally leaves the notion of “multi-hop” abstract, which §4 makes more concrete)
(Global geometric parametric memory) We say that a deep sequence model $f$ has memorized a graph $\mathcal{G}=({V},E)$ geometrically if for any pairs of vertices $u$ and $v$, the logit $f(u)[v]$ reflects some notion of multi-hop closeness in $\mathcal{G}$. At its simplest, this can be abstracted as $f(u)[v]={\bf{\Phi}}_{\tt geom}(u)\cdot{\bf{\Phi}}_{\tt geom}(v)$, where ${\bf{\Phi}}_{\tt geom}$ reflects some graph embedding of the vertices. A fuller definition is in §2.5.
@@
The resolution. We view the two parametric memories of Sections˜2.3 and 2.4 as two competing data structures, both representable by a deep sequence model, each yielding its own learning complexity (much like how a heap or an array would yield different search complexities). Suppose the embeddings of all nodes in path $i$ are clustered tightly around a unique path vector $\mathbf{z}_{i}$. In this data structure, learning the first token given the leaf node is a one-hop task: simply find a one-hop neighbor of the central node that is best-aligned with ${\bf{\Phi}}_{\tt geom}({v}_{\tt{leaf}}^{(i)})$. A similar structure indeed materializes both in our massive graphs and in tiny graphs; this geometry turns out to be fundamental to any gradient-descent-trained deep sequence model and not specific to attention or residual connections:
Figure 4: Evidence of global geometry of Transformer in path-star task. In the heatmap, entry $(i,j)$ is the cosine distance between the leaf embedding of path $i$ (row) and the first-hop embedding of path $j$ (col). The clear diagonal line implies that embeddings within each path are more aligned, reflecting global structure. Analogous plot for Mamba in Fig.˜15
@@
(Evidence of global geometry across various deep sequence models) For our large path-star graphs, the (token) embeddings of the leaf and first node of a path are clustered closer to each other than those of other paths in both the Transformer and Mamba (Figs.˜4 and 15). Similar geometries arise for a variety of small graphs for the Transformer, Mamba and a simple 3-layer neural network model—see visualizations (Fig.˜1 and §C.3: Figs.˜17, 18, 19 and 20) and node-node cosine-similarity heatmaps (§C.3; Figs.˜24 and 23).
@@
While these geometries help make sense of why implicit reasoning succeeds, it also leaves us with a foundational question of why the model prefers one form of memory over the other. We elaborate on this in §3.
2.5 Extended definitions of geometric memory
Below, we provide a more complete and general definition of geometric memory.
Alternative factorizations. The term “geometric memory” can be applied to broader forms of factorized storage where multi-hop information is not present in the logits but can be accessed via a probe. The definition in Section˜2.4, for convenience, implicitly considers a low-rank, positive definite factorization of the adjacency matrix. Two variations in this factorization are possible. One is a full-rank (or non-low-rank) factorization; another factorization is one that is not positive definite but takes the form ${\bf{\Phi}}_{\tt geom}(u)^{T}\mathbf{\Lambda}{\bf{\Phi}}_{\tt geom}(v)$ where $\mathbf{\Lambda}$ is a diagonal matrix of both positive and negative values; such factorizations can arise in multi-layered networks under various settings. However, these factorizations may not yield nice “multi-hop” logit values (as we will see in §4.4). Yet, both these factorizations are geometric in that the nodes are embedded in a highly non-trivial way; one can still extract multi-hop information via a probe trained to detect the global directions of ${\bf{\Phi}}_{\tt geom}$. (The reason why the dominant directions correspond to multi-hop information will be made clear later in §4.2.)
Location of geometry. In the simple settings of this paper, the geometric embeddings ${\bf{\Phi}}_{\tt geom}$ in practice are found in the token embeddings themselves. But more generally, such an embedding may reside in any other layer. We consider even that to be geometric memorization.
Self-edges. The simplistic abstraction in Section˜2.4 may assign the highest output logit to the input vertex itself (i.e., $\arg\max_{v}{\bf{\Phi}}_{\tt geom}(u)\cdot{\bf{\Phi}}_{\tt geom}(v)=u$) regardless of whether a self-loop $(u,u)$ exists in the graph. Neural networks typically seem to handle this well through an auxiliary circuit that suppresses such loops. While we discuss one such mechanism in §4.4, for a tidy discussion, we will ignore it as a pedantic detail.
@@ 3 The memorization puzzle: The emergence of geometric memory is not easy to explain
The reader may offer various plausible explanations for this behavior by borrowing intuition from other well-known phenomena. Some readers may find the rise of a geometry familiar in light of well-known geometries of features in language [44; 140; 121] and arithmetic tasks [108; 129; 126; 64]. Some others may find the phenomenon reminiscent of the competition between memorization and generalization studied in the generalization puzzle [207; 131; 18; 25] in regression and classification settings. These settings suggest that representations arise due to pressures from either the supervision or the architecture or the optimizer; such pressures forbid a lookup. Thus, it may be tempting to cite these well-known learning pressures as plausible explanations for what we observe in our sequence models. However, we design experiments refuting such explanations: we show that a geometric representation arises over the associative lookup even in the absence of such learning pressures. Through this analysis, we pose the memorization puzzle in sequence modeling.
@@ 3.1 Supervisory pressure does not explain geometric memory
Geometric representations are known to arise in various tasks with “multi-hop” objectives, from arithmetic [108; 129; 126; 64] to path-finding [91] to two-hop reasoning [184; 47; 200; 80]. Reasoning supervision has also been shown to act as a regularizer that elicits useful structures in the model’s storage [80]. This trail of findings overwhelmingly directs us to the following guess for why we see a geometry:
(Pressure from global supervision) Perhaps, the model memorizes via a global geometry rather than via local associations due to the supervisory pressure of a multi-hop objective.
This hypothesis however is weakened by the distinctive design of the path-star task: the test paths, having no overlap with the training paths, are untouched by any path-finding supervision, and yet these paths exhibit a global geometry. As an even cleaner refutation of this hypothesis, we ablate the reasoning supervision, and train purely on local supervision i.e., on edge memorization:
(Geometry arises even from memorizing local, atomic co-occurrences) A global geometry emerges even in models trained only on local edge-memorization, as seen in the heatmaps for the large path-star graph (Figs.˜5 and 15(b)). Such models can be subsequently finetuned purely on the hardest-token task and achieve high test accuracy on unseen paths (§D.3). Also, the geometries seen on the small graphs (Fig.˜1 & §C.3) are for various architectures trained only on edge-memorization.
From this, we obtain the following insight: geometries can arise instead of an associative lookup, even when memorizing atomic co-occurrences, in the absence of any pressure to reason or to find high-level, multi-hop structures in data.
Figure 5: Evidence of global geometry of Transformer in edge-memorization task. The geometry observed in Fig.˜4 persists even when we train only on the edges of the path-star task, without path-finding supervision. Mamba shows an even stronger heatmap (Fig.˜15).
3.2 Capacity or optimizer pressures do not explain geometric memory
We may seek yet other explanations out of canonical intuitions in theories of generalization. In regression or classification, conventional wisdom was that models learn representations rather than a lookup when there is an explicit capacity pressure e.g., the architecture or the regularizers forbid a lookup. In modern theories since the popularization of the generalization puzzle [131; 207], renewed wisdom has been that this pressure is implicit: even if the model can express a lookup, gradient descent implicitly prefers solutions that are easier-to-find or are more succinct, thus leading to useful representations. It may seem that one of these theories, when imported to the world of sequence models, can tell us why a geometric representation arises over a naive lookup of atomic associations.